OpenStudy (anonymous):

Transform each polar equation to an equation in rectangular form. Theta = pi/3 Part I: Use the conversion rules to express x and y in terms of r. You should have two equations after this step. Part II. Use the ratio (y/x) to find the slope of the line defined by this equation. Part III: Write the equation of the line in slope-intercept form.

OpenStudy (perl):



OpenStudy (perl):

that is the graph of theta = pi/3 . do you see why? r is allowed to vary so you have polar coordinates (r, pi/3). Now in rectangular coordinates, do you see what you have to do?

OpenStudy (perl):

that is a line through the center with an angle pi/3, and the line has slope tan (pi/3). That means you want y = tan(pi/3) * x

OpenStudy (perl):

another way to do this, in your list of conversion rules you have arctan (y/x) = theta . so you can plug in theta = pi/3 into arctan (y/x) = theta . if you dont reply im leaving the question

OpenStudy (perl):